Hi Guys - Pretty basic algebraic question here however it would be great if you could provide some assistance.
A rectangle is x cm long and y cm wide. There is a frame around the mirrior, which is m cm wide on all four sides.
Questions:
Write a formula for the area of the mirror and frame.
Is your answer a sum or a product, how do you decide?
How can this be answered?
Thanks,
Dom.

Answer by jim_thompson5910(35256)


The length is x. You add m to to each side of this length to get m + x + m = x+2m

So the new length is x+2m cm

Similarly, the old width of y cm becomes y+2m cm

The new area is then...

Area = Length*Width

A = LW

A = (x+2m)(y+2m)

A = xy + 2xm + 2ym + 4m^2

Notice how the second to last line above is a product and the last line is a sum, so either one works.